OpenStudy (anonymous):

5x-10y=-40 solve for y

OpenStudy (anonymous):

\[5x-10y=-40\]we need to get y on one side so whats something you can remove from the side containing y? :)

OpenStudy (anonymous):

5x?

OpenStudy (anonymous):

yes so lets subtract 5x from both sides :)\[~~~5x-10y=-40\]\[\underline{-5x~~~~~~~~~~~~~~-5x}\]\[~~~~~~~-10y=-40-5x\] so now we have \(-10y=-40-5x\)

OpenStudy (anonymous):

\(y\) is still no alone...what can we do?

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

by?

OpenStudy (anonymous):

-5x?

OpenStudy (anonymous):

?...no, try again

OpenStudy (anonymous):

-40

OpenStudy (anonymous):

nope -10

OpenStudy (anonymous):

we need to make the equation look like "\(y=~...\)" so yes we would divide by -10:\[\underline{\cancel{-10}^1y}=\underline{\cancel{-40}^4}~~\cancel{-\underline{~~5~}}^1x\]\[\cancel{-10}^1~~~~~~\cancel{-10}^1~~~~\cancel{-10}^2\]\[{1y\over1}={4\over1}-{1x\over2}\]which is written as:\[y=4+{x\over2}\]

OpenStudy (anonymous):

lets say you have equation \(ax+by=c\), where \(a\) and \(b\) are coefficients and \(c\) is a constant and we are asked to solve for \(y\) to do so we must first get rid of all the terms around the term containing the y, so we would do this:\[ax+by=c\]subtract \(ax\) from both sides:\[by=c-ax\]the \(y\) is still not alone so we will next divide by its coefficient \(b\) to get \(y\) alone:\[by=c-ax\]divide both sides by \(b\):\[y={c\over b}-{ax\over b}\] i hope this better explains things :)

OpenStudy (anonymous):

this is a very helpful video..watch it, you'll probably understand it better :) http://www.youtube.com/watch?v=JPERkt9Ad48

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw.